The height of an object dropped from the roof of an eight story building is modeled by: h(t) -9t^2 + 81, 0
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

The height of an object dropped from the roof of an eight story building is modeled by: h(t) -9t^2 + 81, 0 ======
-------------------
Close.
There's no equal sign.
h(t) = -9t^2 + 81
------------------------
Find h(2)
Find h(0)
---
Then the avg = (h(2) - h(0))/(2-0)

Answer by ikleyn(52778)

.

In your post, EVERYTHING IS WRONG.


One the Earth, where we live, the right equation for the free fall is


    h(t) = -4.9*t^2 + 81,   (if the initial height is 81 meter)


where h(t) is the height in METERS (NOT IN FEET).



In feet,


    h(t) = - 16*t^2 + 81      (if the initial height is 81 feet)


Learn the subject BEFORE trying to compose your own problems.

Learn the subject from these lessons
    - Problem on a projectile moving vertically up and down
    - Problem on an arrow shot vertically upward
    - Problem on a ball thrown vertically up from the top of a tower
    - Problem on a toy rocket launched vertically up from a tall platform

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.
Read them attentively and learn how to solve this type of problems once and for all.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Projectiles launched/thrown and moving vertically up and dawn".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.


//////////////

Do not forget to post your THANKS to me for my teaching.